Project

General

Profile

06-July-2015

Attending:
Enrique, Susan, Bruno, Alex, Chris, Neha, Gavin, Paola, Ken

News (Satish)

  • We will have a production workshop sometime in August to review the first analysis production campaign, and to develop a strategy for streamlining future campaigns.
    The workshop will be two days. To determine the date, Satish will send around a doodle poll.
  • We will be modifying the reco keepup configuration to no longer require presence of the diblock mask, and to produce only blinded cafs. Once we have a new tag with these changes, we will resume reco keepup jobs/
  • As we are now reaching the end of the first-analysis production effort, we will be discontinuing the stand up meetings.

SW Tags (David/Jonathan)

No updates, but we will soon request backport to S15-05-04 series tags to force a crash in the diblock mask for ideal conditions MC. This
is needed for the reconstruction of RHC MC samples.

Simulation (Ruth/Paul)

First analysis simulation jobs are done, except for alternate intensity samples and ND CRY jobs. The former are
Paul's responsibility, but he is till on vacation. Satish will ping Ruth to see if she can run those jobs.

RHC samples have been generated but not yet reconstructed and are waiting for a backport. This has not happened yet because it
was deemed a low-priority item.

Reco/PIDPart (Bruno for Joseph/Enrique/Satish)

  • Data reconstruction is complete. There are a few failures. Most likely these are known BPF crashes. Bruno will confirm.
  • Enrique has just started reco of the the ND Birks-Chou samples. Satish will send instructions on how to submit extra jobs on the
    same SAM project to speed up the process.
  • FD topup MC reconstruction is complete.

LEM (Chris)

Chris is 100% done on with LEM for all inputs and is waiting for ND Birks samples to become available. In the meantime, he has requests to make
decaf files and will proceed with generating them. He will turn off writing of decafs to BlueArc, as this seems to be a major source of FTS backlogs.

Mix/CAF (Bruno/Gavin)

  • FD topup data is mostly done. But Bruno still needs to drain his jobs and make residual restricted CAFs. He also needs decafs to be egnerated, and
    will coordinate with Chris on that.
  • FD birks nonswap and fluxswap are done. The tau files should be done in a few minutes and should be showing up in FTS soon.
  • FD Topup MC is done
  • The FD birksC pid definitions need corrections. Gavin will send around an email update.

Raw2Root Keepup (Paola)

Raw2root running smoothly except for SSL handshake timeout Tue-Wed; An incident was filed, and the problem was resolved.

Calibration Keepup (Paola)

The release for calibration keepup was changed to S15-06-16. Keepup is currently processing with both the new and the old release, using the run date to divide samples and avoid duplication. There were some issues with jobs not running because of the new default memory requirement of 4 GB. The keepup scripts have been updated to now request only 3 GB of memory.

Some jobs at FZU seemed to have a suspiciously long runtime. A GOC ticket was filed, and it turns out the jobs were not running, and FZU was reporting incorrect information. The jobs were killed and resubmitted.